solve |x 2 +4x+3|=x+1 if anyone knows solution please answer asap

solve |x2+4x+3|=x+1 if anyone knows solution please answer asap
 

Grade:12th pass

1 Answers

K VIVEK HRUDAY
4 years ago
|x2+4x+3|=x+1
x2+4x+3=+(or)-[x+1]
This should be split into two cases one with (+)and other with (-)
case1:Lets consider with +
we get x2+4x+3=x+1
            x2+3x+2=0
by splitting into factors we get:
x2+x+2x+2=0
x(x+1)+2(x+1)=0
(x+2)(x+1)=0
solutions are
x=(-1),x=(-2)
by substituting in the equation the solutions satisfies
case2:Let us consider with -
x2+4x+3=(-x)-1
x2+5x+4=0
x2+x+4x+4=0
x(x+1)+4(x+1)=0
(x+4)(x+1)=0
solutions are
x=(-4),x=(-1)
we already got solution as (-1) and we prooved it is right solution(by substituting)
but,if we substitute (-4) it won’t satisfy so number of solutions are two.........
i.e,(-1),(-2).....
HOPE THIS ANSWER HELPS.............